OpenStudy (anonymous):

Give a example of each: 1. Part 1: An absolute value equation with two solutions 2. Part 2: An absolute value equation with one solution 3. Part 3: An absolute value equation with no solutions

OpenStudy (anonymous):

1. Part 1: An absolute value equation with two solutions : abs(x-2) = 6

OpenStudy (anonymous):

2. Part 2: An absolute value equation with one solution : abs(x-1)=0

OpenStudy (anonymous):

3. Part 3: An absolute value equation with no solutions: abs(x-1) =-1

OpenStudy (anonymous):

hope that helps.

OpenStudy (anonymous):

you are literally my lifesaver! thankssss!

OpenStudy (anonymous):

you are welcome